OpenStudy (anonymous):

Write an exponential function y=ab^x for a graph that includes (1,15) and (0,6). a. y=6(2.5)^x b.y=3(5)^x c.y=2.5(6)^x d.y=5(3)^x

myininaya (myininaya):

\[y=a b ^x\] Use the points given to find a and b So we have two equations: \[15=ab^1 \text{ and } 6=ab^0\]

myininaya (myininaya):

\[=>15=ab \text{ and } 6=a(1) \text{ since } b^0=1 \text{ assuming } b \neq 0\]

myininaya (myininaya):

So if a=6 then what is b?

myininaya (myininaya):

And actually you don't really need to know what b equals unless all of the answer choices are wrong lol

OpenStudy (anonymous):

@myininaya \geq \[\geq\] \leq \[\leq\] \Longrightarrow \[\Longrightarrow\] sorry to interrupt O_O

OpenStudy (anonymous):

so it wud be a?

myininaya (myininaya):

@CabDel do you know it is a or do you still have doubts it is a?

OpenStudy (anonymous):

Well since u said A=6 im assuming thats the first number im sure it is.

myininaya (myininaya):

and @kreshnik what is all of that for? What does it have to do with this question? I'm sorry

myininaya (myininaya):

Do you understand how I got a=6?

OpenStudy (anonymous):

Yes kind of.

myininaya (myininaya):

I used the ordered pairs given: \[(x_1,=1,y_1=15) \text{ and } (x_2=0,y_2=6)\] And we have that we know the equation has this form: \[y=ab^x\]

myininaya (myininaya):

\[15=ab^1 \text{ and } 6=ab^0\]

myininaya (myininaya):

I just replaced x with 1 and y with 15 for first ordered pair then I replaced x with 0 and y with 6 for the second ordered pair So I had two equations which I just wrote

myininaya (myininaya):

\[15=ab \text{ and } 6=a\]

myininaya (myininaya):

\[b^0=1 \text{ since } b \neq 0\]

myininaya (myininaya):

=> 15=6b => 15/6=b 15/6 can be written as 2.5 because it is 2.5 lol But yes the answer is a but i just want you to be convinced that it is also a

OpenStudy (anonymous):

Ok thanks im still trying to get used to this chapter its somewhat tough.

myininaya (myininaya):

You can do it! OpenStudy is here to help and root for ya!

OpenStudy (anonymous):

Yeah hah thanks!. Anyway i gtg thanks for the help